A man is paid Php300 for each day he works and forfeits Php80 for each day he is idle. At the end of 20 days he nets Php4,500. How many days did he work?
Answer by stanbon(75887)

A man is paid Php300 for each day he works and forfeits Php80 for each day he is idle. At the end of 20 days he nets Php4,500. How many days did he work?
------------
Let # of days he worked be "x"; His pay is 300x
-------------
Then "20-x" is # of days he did not work: His loss = 80(20-x)= 1600-80x
------------------------
EQUATION:
Gain - Loss = 4500
300x -(1600-80x) = 4500
380x = 6100
x = 16.0526 days ( which doesn't make any sense)
He worked 16.0526 days
==========================
Cheers,
Stan H.